Owl-Eye Radiologists (OR) does various types of diagnostic imaging. Radiologists perform tests using sophisticated equipment. OR's management wants to compute the costs of performing tests for two different types of patients: those who are hospitalized (including those in emergency rooms) and those who are not hospitalized but are referred by physicians. The data for June for the two categories of patients follow.

Hospital Patients Other Patients Total
Units (i.e., procedures) produced 670 800 1,470
Equipment-hours used 250 150 400
Direct labor-hours 510 190 700
Direct labor costs $ 40,800 $ 11,400 $ 52,200
Overhead costs $ 51,700

The accountant first assigns overhead costs to two pools: overhead related to equipment-hours and overhead related to labor-hours. The analysis of overhead accounts by the cost accountant follows.

Account Amount Related to:
Utilities $ 4,840 Equipment-hours
Supplies 12,580 Labor-hours
Indirect labor and supervision 20,320 Labor-hours
Equipment depreciation and maintenance 10,460 Equipment-hours
Miscellaneous 3,500 Equipment-hours


Required:

a. Compute the predetermined overhead rates assuming that Owl-Eye Radiologists uses equipment-hours to allocate equipment-related overhead costs and labor-hours to allocate labor-related overhead costs.

b. Compute the total costs of production and the cost per unit for each of the two types of patients undergoing tests in June.
Answer #1

a.

Equipment related overhead costs = Utilities + Equipment depreciation and maintenance + Miscellaneous

= 4,840+10,460+3,500

= $18,800

Labor related overhead costs = Supplies +  Indirect labor and supervision

= 12,580+20,320

= $32,900

Predetermined overhead rate for equipment related overheads = Equipment related overhead costs/ Equipment-hours used

= 18,800/400

= $47 per equipment hour

Predetermined overhead rate for labor = Labor related overhead costs / Direct labor-hours

= 32,900/700

= $47 per direct labor hour

b.

Hospital Patient Other Patient
Direct labor cost 40,800 11,400
Equipment related overheads 250 x 47 = 11,750 150 x 47 = 7,050
Labor related overheads 510 x 47 = 23,970 190 x 47 = 8,930
Total cost 76,520 27,380
Number of units 670 800
Cost per unit (Total cost/ Number of units) 114.21 34.23
Hospital Patient Other Patient
Total cost $76,520 $27,380
Cost per unit (Total cost/ Number of units) $114.21 $34.23

Kindly comment if you need further assistance. Thanks‼!